# The Dark Side of Rich Rewards: Understanding and Mitigating Noise in VLM Rewards

### v4

@@ 5 False Positives From Cosine Similarity
The Issue of State Entanglement State entanglement refers to the issue where the cosine similarity metric erroneously pays more attention to lexical-level similarity while lacking comprehension of the underlying state transitions. Consequently, rewards are given to trajectory-instruction pairs that are cosine similar in embedding space but in fact result in distinct state transitions. For instance, consider the significant contrast between “take the red pill” and “take the blue pill”. Despite their lexical similarity, they lead to vastly different states. However, the cosine similarity metric may represent them as similar due to the shared words, disregarding the critical difference in state outcomes. Understanding state transitions is crucial in sequential decision-making scenarios. Otherwise, rewards may be given to trajectories that lead to unintended states, potentially prolonging the path to the goal state by necessitating corrective actions or re-attempts.
@@
The Issue of Composition Insensitivity Composition insensitivity in cosine similarity metrics gives rise to two issues: (1) rewarding incomplete task execution – cosine similarity may incorrectly reward partial task completion, as even incomplete trajectories can receive high similarity score in the embedding space. For instance, in a task to “pick up a tool, then fix a machine,” the model might prematurely reward the agent for merely picking up the tool, neglecting the crucial repair action. We also observed this phenomenon particularly in the Montezuma environment, where RL agents tend to hack the reward system by focusing on the easiest actions that yield rewards (e.g., moving towards a direction) rather than executing more complex, timely actions. Eventually, this leads to an overestimation of the agent’s progress towards the ultimate goal. (2) insensitivity to the ordering of execution – cosine similarity often fails to adequately penalize incorrect execution sequences. In a safety protocol requiring an agent to “turn off the machinery, then perform maintenance,” the metric might assign high rewards based merely on the presence of relevant actions, disregarding their order. In contrast to some advancements in language models, compact visual and sentence embeddings from multimodal VLMs remain largely insensitive to sequential information (Pham et al., 2021). When the task is order-sensitive, executing actions in the wrong sequence prolongs the path towards the goal state, as agents need to re-attempt the correct order.
Figure 1 illustrates various scenarios where false positive rewards are erroneously assigned. To empirically demonstrate the issue, Section 5.1 presents experiments on these issues and their impact on agent learning in sparse reward environments.
@@
5.1 Experiments on Reward Noise Impact
Our experiments test the following hypothesis: (H1) The two issues of state entanglement and composition insensitivity exist; (H2) false positive rewards are prevalent during training; (H3) VLM reward models lacking noise handling mechanisms underperform against intrinsic reward models in sparse reward environments; (H4) false negatives may not be as harmful as false positives.
Setup. We evaluate these hypotheses through various challenging sparse-reward environments: (1) Crafter, an open-ended 2D Minecraft (Hafner, 2021); (2) Montezuma, a classic hard adventure game in Atari (Bellemare et al., 2013); and (3) Minigrid ‘Go To Seq’, a hard task involving long-horizon navigation and object interactions (Chevalier-Boisvert et al., 2018). A Markovian and a Non-Markovian reward model were tested: (1) Pixl2R by Goyal et al. (2020), which uses only the current video frame to determine if the goal state specified in the instruction has been reached; and (2) ELLM-, a variant of ELLM by Du et al. (2023). Unlike ELLM, which queries instructions from LLMs in real-time, ELLM- directly uses preset expert instructions and compares them with the transition differences of the agent’s trajectory. The VLM backbones used are: (1) CLIP (Radford et al., 2021), pretrained by image-text pairs; and (2) X-CLIP (Ma et al., 2022), pretrained by video-text pairs. To ensure high-quality finetuning data, we used internal information from the game engine to annotate expert trajectories from expert agents. To demonstrate how noisy reward signals hinder learning, we selected a strong intrinsic reward model DEIR (Zhang et al., 2021) for comparison. It provides auxiliary rewards based on observation novelty to encourage exploration. See Appendix A.4 for detailed implementation of the experiments.
Evaluation Metric. We adopted the score metric from the Crafter benchmark (Hafner, 2021) for performance evaluation, as it effectively measures consistent performance across multiple subtasks in sparse reward environments. Unlike the maximum total rewards metric, which does not adequately reflect consistent performance, the score metric offers a more reliable indicator of learning progress. See Appendix A.5 for its formal definition.
Reward Noise Issue. To investigate H1, we evaluated the models’ sensitivity by examining how cosine similarity scores change for manipulated trajectory-instruction pairs. The state entanglement test involved reversing trajectories and negating instructions (i.e., “do not do $l_{k}$”). The composition insensitivity test examined concatenated pairs of trajectory-instruction data. Given $(\tau_{1},l_{1})$ and $(\tau_{2},l_{2})$, we create a concatenated pair $(\tau_{1}+\tau_{2},l_{1}+l_{2})$. We then test two types of manipulations – (1) swapping the order within one modality: e.g., $(\tau_{2}+\tau_{1},l_{1}+l_{2})$; and (2) truncating one modality: e.g., $(\tau_{1},l_{1}+l_{2})$. Overall, three types are evaluated: (1) matched pairs; (2) not-matched pairs and (3) manipulated pairs which are derived from matched pairs by polluting either the trajectory or the instruction. Our results reveal a critical flaw in the reward model: despite manipulated pairs that fundamentally fail to fulfill the instruction, the model paradoxically assigns high similarity scores (see Figure 4 for overall results and also Appendix A.6 for individual environments). It’s worth noting that the poor performance in the negation case aligns with broader challenges in natural language processing. Recent studies (Hossain et al., 2022; Truong et al., 2023) have highlighted that negation is central to language understanding but is not properly captured by modern language models. This limitation extends to VLMs and directly leads to false positive rewards.
@@ 5.1 Experiments on Reward Noise Impact
Pixl2R VLM $0.142$ $12.4$ $9.40$ $\mathbf{-49\%}$
ELLM- VLM $0.150$ $19.4$ $10.8$ $-41\%$
Pixl2R + DEIR VLM + intr. $0.176$ $17.3$ $10.4$ $-38\%$
ELLM- + DEIR VLM + intr. $0.178$ $30.9$ $11.8$ $-27\%$
Prevalence of False Positives. To address H2, we analyzed reward distribution heatmap from VLM-based reward models during training. The heatmap revealed a concerning trend: RL agents engage in reward hacking, receiving rewards across vast areas of the environment rather than just at goal states. For instance, in Montezuma where the goal is to grab the key and escape the room, we observed that agents received rewards even for falling off cliffs, which undoubtedly contribute to false positive rewards. For environments without fixed camera views, we calculated the step offset between the current rewarded state and the actual goal state. A positive offset indicates a false positive reward, as the reward was given before reaching the goal. Conversely, a negative offset indicates a false negative, where the agent reached the goal but the reward model failed to acknowledge it (see Figure 6). Interestingly, besides positive offsets, we observed a large amount of negative offsets in Minigrid environments. We attribute this to Minigrid’s abstract shape-based visual representations, which fall outside the VLM’s training distribution.
Impact on Learning. We trained agents using learned VLM reward models and compared their learning efficacy against intrinsic reward models. As shown in Table 1, our results confirmed H3: instruction-following RL agents using learned VLM reward models without noise handling consistently underperform compared to DEIR, the intrinsic reward-based RL agent. To investigate the impact of false negatives versus false positives (H4), we designed an oracle Pixl2R model with two variants: a false negative model and a false positive model. The false negative model only rewards the agent for reaching subgoal states described in the instruction, with a probability of xthat some rewarding states in the map are removed. In contrast, the false positive model rewards the agent for reaching every subgoal, but also introduces a small (0.1) one-off reward for certain locations, covering xof the map. The results indicate that false negatives were less detrimental to agent performance than false positives (see Figure 4). This performance difference can be explained through our theoretical analysis in Appendix A.2.2.
@@ 6 Addressing the False Positive Issue
Our proposed solution is not specific to cosine similarity’s approximation error, which serves as a case study. Rather, it’s a broad strategy for mitigating false positive rewards across various reward models, including noisy VLM reward sources.
6.1 Binary Signal and Conformal Prediction
@@
Our experiments have demonstrated that false positives are more detrimental to learning than false negatives. Based on these findings, we propose a reward function that issues a one-time binary reward only when the similarity between the agent’s current trajectory and the instruction exceeds a high confidence threshold. This approach contrasts with previous methods, which provide continuous rewards whenever the reward score exceeds a predefined threshold, and continue to do so until reaching a maximum cap. Our method, however, delivers this reward only once. This approach minimizes the likelihood of accumulating false positive rewards while maintaining adherence to Assumption 4.3.
@@
To achieve this, we introduce a thresholding mechanism using a calibration set of true positive trajectory-instruction pairs. This threshold, denoted as $\hat{q}$, is set to the empirical quantile of cosine similarity scores at the significance level $1-\alpha$. Pairs whose similarity scores fall below this threshold $\hat{q}$ receive no reward. Conversely, pairs exceeding $\hat{q}$ receive a one-time $+1$ reward, i.e., $r^{v}_{\textsc{Bi}}(\tau,l_{k})=\mathbf{1}_{\{p(l_{k}\mid\tau)\geq\hat{q}\}}$. This thresholding approach leverages statistical properties studied by Sadinle et al. (2019), which ensures a high probability (at least $1-\alpha$) that true positive pairs are recognized while minimizing frequency of false positives errors. See Appendix A.4.2 for the pseudo-code of the instruction-following RL algorithm and Appendix A.8 for detailed threshold calculation.
6.2 Mutual Information Maximization
Intuitively, when we observe rewards coming from a particular signal source too frequently, we tend to downplay the significance of that signal to avoid over-reliance. This intuition is effectively captured by incorporating a mutual information maximization term into the reward function. Specifically, the updated reward function $r^{v}_{\textsc{MI}}(\tau,l_{k})$ measures the mutual information between the agent’s trajectory and the instruction. Mathematically, it can be expressed as:
$=\mathbb{E}_{\tau\sim\pi_{\theta},l_{k}\sim L}[\log p(l_{k}\mid\tau)-\log p(l_{k})]$
where $\tau=\langle s_{t-W},a_{t-W},\ldots,s_{t}\rangle$ is the agent’s trajectory up to current time step $t$, and $W$ is the memory size of the agent for its past trajectory. $p(l_{k}\mid\tau)$ comes from the similarity score provided by VLMs, referring to the likelihood of the instruction $l_{k}$ being fulfilled by a trajectory $\tau$. $p(l_{k})$ is overall likelihood of encountering the instruction $l_{k}$ in the learning environment. Therefore, the second term in the equation serves as a regularization term that downplays the significance of the reward signal when it is too frequent. For instance, if a VLM frequently detects that the agent’s actions are fulfilling the “climbing the ladder” instruction, even when the agent is performing unrelated tasks, any reward signal from this instruction will be downplayed. $p(l_{k})$ is calculated as follows:
$p(l_{k})=\mathbb{E}_{\tau\sim\pi_{\theta_{-1}}}\left[\left(\sum\nolimits_{t=1}^{T_{\tau}}\mathbf{1}_{\{p(l_{k}\mid\tau_{t})\geq\hat{q}\}}\right)/T_{\tau}\right]$
Here, $\tau_{t}$ is the agent’s trajectory up to time $t$, and if the VLM deems the trajectory as fulfilling the instruction (i.e., $p(l_{k}\mid\tau_{t})\geq\hat{q}$), we increment the count. Dividing the count by the total trajectory length $T_{\tau}$ gives the empirical frequency of the instruction being fulfilled. The subscript $\theta_{-1}$ in $\pi_{\theta_{-1}}$ indicates that the trajectories are sourced from rollouts in the previous policy iteration, acknowledging the impracticality of real-time computation of $p(l_{k})$ during an ongoing episode.
@@ 6.2 Mutual Information Maximization
Table 2: Model score across various environments. $\star$ is the baseline agents with a learned VLM-based reward model to compare with. BiMI significantly improves performance in Montezuma and Minigrid, while showing mixed results in Crafter due to task-specific characteristics
@@
Pixl2R + Bi $0.137\pm 0.009$ $-3.5\%$ $31.2\pm 2.04$ $+151\%$ $10.7\pm 0.784$ $+14\%$
Pixl2R + BiMI $0.162\pm 0.022$ $+14\%$ $37.5\pm 7.83$ $+199\%$ $7.95\pm 0.351$ $-15\%$
Pixl2R + DEIR $0.176\pm 0.009$ $+23\%$ $17.3\pm 0.51$ $+39\%$ $10.4\pm 1.015$ $+10\%$
Pixl2R + BiMI + DEIR $0.267\pm 0.016$ $+88\%$ $57.7\pm 2.15$ $+364\%$ $11.0\pm 0.190$ $+17\%$
@@
ELLM- + Bi $0.151\pm 0.016$ $+0.6\%$ $29.7\pm 1.29$ $+53\%$ $11.1\pm 0.601$ $+3.2\%$
ELLM- + BiMI $0.156\pm 0.014$ $+4.0\%$ $33.6\pm 3.99$ $+74\%$ $9.42\pm 0.267$ $-12\%$
ELLM + DEIR $0.178\pm 0.029$ $+20\%$ $30.9\pm 3.50$ $+59\%$ $11.8\pm 1.152$ $+9.5\%$
ELLM- + BiMI + DEIR $0.279\pm 0.078$ $+86\%$ $56.2\pm 6.19$ $+190\%$ $13.1\pm 0.393$ $+22\%$
@@
It’s important to note that the BiMI approach primarily mitigates false positives (FP) rather than false negatives (FN). Both Bi and MI aim to reduce the likelihood of rewarding unintended trajectories, thus addressing the FP issue. While this conservative approach may increase false negatives by pruning some correct trajectories, this trade-off is beneficial, as demonstrated by both our empirical and theoretical results.
7 Experiments
We continue to evaluate using Markovian Pixl2R and Non-Markovian ELLM- and their BiMI-enhanced counterparts, while also exploring potential synergies with the intrinsic reward model DEIR. We follow the same experimental setup as in Section 5.1, with additional details in Appendix A.9.
7.1 Main Results

### v5

@@ 5 False Positives From Cosine Similarity
Figure 2: Schematic diagram of false positives in embedding space.
@@
The Issue of State Entanglement State entanglement refers to the issue where the cosine similarity metric erroneously pays more attention to lexical-level similarity while lacking comprehension of the underlying state transitions. Consequently, rewards are given to trajectory-instruction pairs that are cosine similar in embedding space but in fact result in distinct state transitions. For instance, consider the significant contrast between “take the red pill” and “take the blue pill”. Despite their lexical similarity, they lead to different states. However, the cosine similarity metric may represent them as similar due to the shared words, disregarding the critical difference in state outcomes (see Figure 2 top left). Understanding state transitions is crucial in sequential decision-making scenarios. Otherwise, rewards may be given to trajectories that lead to unintended states, potentially prolonging the path to the goal state by necessitating corrective actions or re-attempts.
@@
The Issue of Composition Insensitivity One might wonder why composition insensitivity persists despite using a pointer mechanism (Appendix C) that enforces sequential instruction order. The issue arises because sentences $l$ in the instruction sequence $L$ are not always atomic: a single instruction like “tidy the room before leaving” may implicitly encode multiple sub-tasks. While the pointer ensures progress through high-level steps, it does not help the VLM parse the internal structure of complex, non-atomic instructions. As a result, “composition insensitivity” emerges, misaligning the agent’s trajectory and undermining reward accuracy. Unfortunately, even state-of-the-art natural language understanding models struggle with decomposing atomic sentences without losing critical sequential or contextual information (Dziri et al. 2024). Under this circumstance, composition insensitivity in cosine similarity metrics gives rise to two issues: (1) rewarding incomplete task execution – cosine similarity may incorrectly reward incomplete task execution by giving high scores even when critical steps are missing. We observed this phenomenon particularly in the Montezuma environment, where RL agents tend to hack the reward system by focusing on the easiest actions that yield rewards (e.g., moving towards a direction) rather than executing more complex, timely actions. Eventually, this leads to an overestimation of the agent’s progress towards the ultimate goal (see Figure 2 bottom left). (2) insensitivity to the ordering of execution – VLM models often fails to adequately penalize incorrect execution sequences, rather, it assigns high rewards based merely on the presence of relevant actions, disregarding their order (see Figure 2 top right). In contrast to some advancements in language models, compact visual and sentence embeddings from multimodal VLMs remain largely insensitive to sequential information (Pham et al. 2020). When the task is order-sensitive, executing actions in the wrong sequence prolongs the path towards the goal state, as agents need to re-attempt the correct order — yet the VLM’s reward may still incentivize the agent to continue with the incorrect sequence. We empirically demonstrate these issues in Section 6, showing their substantial impact on agent learning in sparse reward environments.
6 Experiments on Reward Noise Impact
Our experiments test the following hypothesis: (H1) The two issues of state entanglement and composition insensitivity exist, (H2) false positive rewards are prevalent during training, (H3) VLM reward models lacking noise handling mechanisms underperform against intrinsic reward models in sparse reward environments, (H4) false negatives may not be as harmful as false positives.
@@ 6 Experiments on Reward Noise Impact
Experimental Setup We evaluate these hypotheses through various challenging sparse-reward environments: (1) Crafter, an open-ended 2D Minecraft (Hafner 2022); (2) Montezuma, a classic hard adventure game in Atari (Bellemare et al. 2013); and (3) Minigrid ‘Go To Seq’, a hard task involving long-horizon navigation and object interactions (Chevalier-Boisvert et al. 2018). Details on why these environments are preferred over other test benchmarks, along with the evaluation metrics, are provided in Appendix A. A Markovian and a Non-Markovian reward model were tested: (1) Pixl2R by Goyal, Niekum, and Mooney (2020), which uses only the current video frame to determine if the goal state specified in the instruction has been reached; and (2) ELLM-, a variant of ELLM by Du et al. (2023) that directly uses preset expert instructions and compares them with the transition differences of the agent’s trajectory. The VLM backbones used are: (1) CLIP (Radford et al. 2021), pretrained by image-text pairs; and (2) X-CLIP (Ma et al. 2022), pretrained by video-text pairs. To ensure high-quality finetuning data, we used internal information from the game engine to annotate expert trajectories from expert agents. To demonstrate how noisy reward signals hinder learning, we selected a strong intrinsic reward model DEIR (Wan et al. 2023) for comparison. It provides auxiliary rewards based on observation novelty to encourage exploration. The detailed VLM-based reward model procedure, finetuning process and training hyperparameters are in Appendix H.
Reward Noise Issue To investigate H1 — the existence of two issues, state entanglement and composition insensitivity, in VLM-based reward models — we evaluated the models’ sensitivity by examining how cosine similarity scores change for manipulated trajectory-instruction pairs. These manipulations, designed to probe to robustness against noise, included the following:
  1. 1.
  2. 2.
  3. 3.
    Concatenation and Order Swapping: Given two trajectory-instruction pairs $(\tau_{1},l_{1})$ and $(\tau_{2},l_{2})$, we created concatenated pairs and then swapped the order in one modality. For example: a) Original concatenation: $(\tau_{1}+\tau_{2},l_{1}+l_{2})$; b) Swapped trajectory: $(\tau_{2}+\tau_{1},l_{1}+l_{2})$; c) Swapped instruction: $(\tau_{1}+\tau_{2},l_{2}+l_{1})$. This tests the model’s sensitivity to the order of components in multi-step tasks.
  4. 4.
    Concatenation with Partial Content: We concatenated pairs but truncated one modality. For instance: a) Truncated trajectory: $(\tau_{1},l_{1}+l_{2})$; b) Truncated instruction: $(\tau_{1}+\tau_{2},l_{1})$. This assesses the model’s ability to detect partial mismatches in longer sequences.
Our results reveal a critical flaw in the reward model: for manipulated pairs that fundamentally fail to fulfill the instruction, the model paradoxically assigns high similarity scores (see Figure 4 for overall results and also Appendix I for individual environments). Note that the poor performance in the negation case aligns with broader challenges in natural language processing. Recent studies (Hossain et al. 2022; Truong et al. 2023) have highlighted that negation is central to language understanding but is not properly captured by modern language models. This limitation extends to VLMs and directly leads to false positive rewards.
Prevalence of False Positives To address H2, we analyzed reward distribution heatmap from VLM-based reward models during training. The heatmap (Figure 6 left) revealed a concerning trend: RL agents engage in reward hacking, receiving rewards across vast areas of the environment rather than just at goal states. For instance, in Montezuma where the goal is to grab the key and escape the room, we observed that agents received rewards even for falling off cliffs, which undoubtedly contribute to false positive rewards. For environments without fixed camera views, we calculated the step offset between the current rewarded state and the actual goal state. A positive offset indicates a false positive reward, as the reward was given before reaching the goal. Conversely, a negative offset indicates a false negative, where the agent reached the goal but the reward model failed to acknowledge it (see Figure 6). Notably, besides positive offsets, we observed a large amount of negative offsets in Minigrid environments. We attribute this to Minigrid’s abstract shape-based visual representations, which fall outside the VLM’s training distribution.
Impact on Learning We trained agents using learned VLM reward models and compared
@@
[…]
their learning efficacy against intrinsic reward models. As shown in Table 1, our results confirmed H3: instruction-following RL agents using learned VLM reward models without noise handling consistently underperform compared to DEIR, the intrinsic reward-based RL agent.
In our efforts to assess the impact of false positive rewards from auxiliary reward model without the interference of other factors such as domain shift, poor data quality, and errors from other issues such as the choice of multimodal architectures, we devised a simulated reward model that access to internal state information from the game engine. Experiments with these simulated models provide initial evidence of the differential impact of false negatives versus false positives on training outcomes, as posited by H4, suggesting that false positives — particularly those tied to temporal insensitivity — more severely degrade final scores. Due to page limits, we move the details to Appendix J.
To compare the effects of reward noise, we designed an oracle Pixl2R model with two variants: one simulating false negatives by randomly removing $x\%$ of true subgoal rewards, and one simulating false positives by adding small one-off rewards (0.1) to $x\%$ of irrelevant states. This design reflects a realistic failure mode of VLMs, where false positive rewards — despite being only 1/10 the magnitude of true subgoal rewards and sometimes located far from the optimal trajectory — can still be exploited by the agent through training iterations, a phenomenon known as reward hacking. In contrast, false negatives merely reduce the frequency of positive feedback without misleading the agent. The results indicate that false negatives were less detrimental to agent performance than false positives (see Figure 4). This aligns with our theoretical analysis in Theorem 4.10.
@@ 7 Addressing the False Positive Issue
Binary Signal and Conformal Prediction  Having established the detrimental effects of false positive rewards, our solution is to strategically manage the trade-off: reducing false positives, even at the cost of a slight increase in false negatives. First, we propose a reward function that issues a one-time binary reward only when the similarity between the agent’s current trajectory and the instruction exceeds a high confidence threshold. This approach contrasts with previous methods, which provide continuous rewards whenever the reward score exceeds a predefined threshold, and continue to do so until reaching a maximum cap. Our method, however, delivers this reward only once. This approach minimizes the likelihood of accumulating false positive rewards while maintaining adherence to Proposition 4.4.
@@
To achieve this, we introduce a thresholding mechanism using a calibration set of true positive trajectory-instruction pairs. This threshold, denoted as $\hat{q}$, is set to the empirical quantile of cosine similarity scores at the significance level $1-\alpha$. Pairs whose similarity scores fall below this threshold $\hat{q}$ receive no reward. Conversely, pairs exceeding $\hat{q}$ receive a one-time +1+1 reward:
$r^{v}_{\textsc{Bi}}(\tau,l_{k})=\mathbf{1}_{\{p(l_{k}\mid\tau)\geq\hat{q}\}}$ (1)
It statistically guarantees a high probability (at least $1-\alpha$) that the true positive pairs are recognized (i.e., above the threshold) while minimizing the average number of mistakes predicting false positives (Sadinle et al. 2019). The threshold calculation is detailed in Algorithm 2.
@@
Table 2: Model score across various environments. $\star$ is the baseline agents with a learned VLM-based reward model to compare with. BiMI significantly improves performance in Montezuma and Minigrid, while showing mixed results in Crafter
@@
[…]
Pixl2R + DEIR $0.176\pm 0.009$ $23\%$ $17.3\pm 0.51$ $39\%$ $4\pm 1.015 +10+10\%$
[…]
@@
[…]
ELLM + DEIR $0.178\pm 0.029$ $20\%$ $30.9\pm 3.50$ $59\%$ $11.8\pm 1.152$ $9.5\%$
[…]
@@
Mutual Information Maximization Intuitively, when we observe rewards coming from a particular signal source too frequently, we tend to downplay the significance of that signal to avoid over-reliance. This intuition is effectively captured by incorporating a mutual information maximization term into the reward function. Specifically, the updated reward function $r^{v}_{\textsc{MI}}(\tau,l_{k})$ measures the mutual information between the agent’s trajectory and the instruction. Formally, it can be expressed as:
$=\mathbb{E}_{\tau\sim\pi_{\theta},l_{k}\sim L}[\log p(l_{k}\mid\tau)-\log p(l_{k})]$ (2)
where $\tau=\langle s_{t-W},a_{t-W},\ldots,s_{t}\rangle$ is the agent’s trajectory up to current time step $t$, and $W$ is the memory size of the agent for its past trajectory. $p(l_{k}\mid\tau)$ comes from the similarity score provided by VLMs, referring to the likelihood of the instruction $l_{k}$ being fulfilled by trajectory $\tau$. $p(l_{k})$ is overall likelihood of encountering the instruction $l_{k}$ in the learning environment. Therefore, the second term in the equation serves as a regularization term that downplays the significance of the reward signal when it is too frequent. For instance, if a VLM frequently detects that the agent’s actions are fulfilling the “climbing the ladder” instruction, even when the agent is performing unrelated tasks, any reward signal from this instruction will be downplayed.
We treat the set of instructions $\{l_{k}\}$ as a finite discrete support and define an empirical frequency:
$s(l_{k})\;=\;\mathbb{E}_{\tau\sim\pi_{\theta_{-1}}}\!\Bigl[\tfrac{1}{T_{\tau}}\sum_{t=1}^{T_{\tau}}\mathbf{1}\bigl\{p(l_{k}\mid\tau_{t})\geq\hat{q}\bigr\}\Bigr$], (3)
where $\tau_{t}$ is the agent’s trajectory up to time $t$, and we count one whenever the VLM scores $\tau_{t}$ as fulfilling $l_{k}$ over the total trajectory length $T_{\tau}$. We then normalize to obtain a probability mass function (PMF) over instructions: $p(l_{k})\;=\;\frac{s(l_{k})}{\sum_{j}s(l_{j})}$, and $\sum_{j}s(l_{j})=1$. Because $\{l_{k}\}$ is discrete, $p(l_{k})$ is a valid PMF (not a continuous density) and can be plugged directly into the usual discrete mutual information in Section 7. The subscript $\theta_{-1}$ in $\pi_{\theta_{-1}}$ indicates that the trajectories are sourced from rollouts in the previous policy iteration, acknowledging the impracticality of real-time computation of $p(l_{k})$ during an ongoing episode.
@@
Note that the BiMI approach primarily mitigates false positives (FPs) rather than false negatives (FN). Both Bi and MI aim to reduce the likelihood of rewarding unintended trajectories, thus addressing the FP issue. While this conservative approach may increase false negatives by pruning some correct trajectories, this trade-off is beneficial, as demonstrated by both our empirical and theoretical results.
8 Experiments and Results
We follow the same experimental setup as in Section 6 to ensure a fair comparison. Furthermore, we set confidence level for empirical quantile calculation to be $1-\alpha=0.9$. We adhered to the standard requirement of limiting the training budget to 1 million frames (Hafner 2022).